OpenStudy (anonymous):

The Doe family is ready to fill their new swimming pool. It can be filled in 12 hours if they use their own water hose, and in 30 hours if they use Mr. Jones', their neighbor's water hose. How long will the Doe's take to fill their pool if the neighbor's hose is used along with their own?

OpenStudy (irishboy123):

Use algebra. Let V = pool vol Let d = flow rate for doe's hose (volume per hour) Let j = same but for the jones' You can then say V = 12d And V = 30j You then need to calc V/(d+j) Well you know that 12d = 30j,..., so finish it off!

OpenStudy (zpupster):

or 1/12 +1/30 = 1/x